OpenStudy (anonymous):

simplify (v^-7)4 write without the negative sgin

OpenStudy (anonymous):

You mean \[(v^{-7})^4 ?\]

OpenStudy (anonymous):

1/v^28

OpenStudy (anonymous):

\[(v^{-7})4\]

OpenStudy (anonymous):

so \[4 \times v^{-7}? \] Only way to simplify is \[4 \times v^{-7}=\frac{4}{v^7}\]

OpenStudy (anonymous):

yes!!!!!! thanks :D

OpenStudy (anonymous):

v-4/5=4 1/3